a 10 pound mixture of peanuts and cashews sells for $5.32 per pound. the price of peanuts is $3.60 per pound and the price of cashews is $7.90 per pound. how many pounds of each type are in the mixture?
Answer by checkley71(8403)

7.9*X+3.6(10-X)=5.32*10
7.9X+36-3.6X=53.2
4.3X=53.2-36
4.3X=17.2
X=17.2/4.3
X=4 ANSWER FOR THE POUNDS OF CASHEWS USED.
10-4=6 ANSWER FOR THE POUNDS OF PEANUTS USED.
PROOF:
7.9*4+3.6*6=53.2
31.6+21.6=53.2
53.2=53.2